# SynCo: Synthetic Hard Negatives in Contrastive Learning for Better Unsupervised Visual Representations

### v1

@@ 4.1 Hard Negatives Generation
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\text{sim}(\cdot,\cdot)$ is the similarity function and $\delta$ controls the perturbation magnitude. We approximate the gradients by calculating the difference in similarity scores with small perturbations applied. The perturbed embeddings are then normalized and added to the negative logits. Perturbation modifies the embeddings of the hardest negative features based on the gradient of the similarity function, creating synthetic negatives that are slightly adjusted to be more challenging for the model.
Adversarial hard negatives (type 6). For each query $\mathbf{q}$, we propose to generate $N_{6}$ hard negative features by applying adversarial perturbations to the hardest negative features to maximize their similarity to the query embeddings. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{6}=\{\mathbf{h}_{1}^{6},\mathbf{h}_{2}^{6},\ldots,\mathbf{h}_{N_{6}}^{6}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{6}\in H^{6}$ would be given by:
@@ 4.2 Discussion on Synthetic Hard Negatives
Intuition of synthetic hard negatives. Each SynCo strategy enhances model generalization through challenging contrasts. Type 1 interpolates between query and hard negatives, increasing sample diversity throughout training. Type 2 extrapolates beyond the query, pushing embedding space boundaries and improving robustness to difficult contrasts. Type 3 combines pairs of hard negatives, encouraging more generalized and robust feature learning. Type 4 injects Gaussian noise, promoting invariance to minor feature fluctuations and enhancing generalization. Type 5 modifies embeddings based on similarity gradients, refining discriminatory power by directing the model towards harder negatives. Type 6 applies adversarial perturbations, creating the most challenging contrasts to distinguish deceptively similar samples. A toy example of the proposed synthetic hard negative generation is presented in Figure 4. The latter figure visualizes the different types of synthetic hard negatives generated from a query point and three original negative points in a 2D embedding space. Dotted lines connect the query to extrapolated negatives and between pairs of mixup negatives, illustrating the relationships used to generate these synthetic points.
@@
Figure 4: A toy example illustrating the six types of synthetic hard negatives generated by SynCo from a query point and three original negative points in a 2D embedding space.
@@
Is the proxy task more difficult? Figure 3 illustrates the proxy task performance for various configurations of SynCo. We observe that incorporating synthetic negatives leads to faster learning and improved performance. Each type of synthetic negative accelerates learning compared to the MoCo-v2 baseline, with the full SynCo configuration showing the most significant improvement (see Figure 6(a)) and the lowest final proxy task performance. This indicates that SynCo presents the most challenging proxy task. The lower performance when including synthetic negatives suggests that $\max\ell(\mathbf{h}_{k}^{i})>\max\ell(\mathbf{n}_{j})$, $\mathbf{h}_{k}^{i}\in H^{i}$, $\mathbf{n}_{j}\in\tilde{\mathcal{Q}}_{N}$. Through SynCo, we modulate proxy task difficulty via synthetic negatives, pushing the model to learn more robust features.
@@
Evaluating the usage of the embedding space. Wang and Isola [38] presented two losses/metrics for assessing contrastive learning representations. The first metric assesses alignment by measuring the absolute distance between representations that share the same label on a hypersphere. The second metric evaluates uniformity by calculating the logarithm of the average pairwise Gaussian potential among all embeddings. In Figure 5, we present these metrics for various models, based on re-implemented methods and experiments conducted using features from all images in the ImageNet-100 validation set. Our findings demonstrate that SynCo significantly enhances the uniformity of representations compared to MoCo-v2 and MoCHI, indicating its superior ability to utilize the embedding space in the proxy task. Furthermore, the incorporation of synthetic negatives (types 1 to 6) leads to improved alignment. These results suggest that SynCo’s approach to synthetic negative generation and contrastive learning yields more effective and well-distributed feature representations.
@@
Figure 5: Performance comparison of MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of alignment and uniformity metrics. The x/y axes correspond to $-\mathcal{L}_{\text{uniform}}$ and $-\mathcal{L}_{\text{align}}$, respectively. The best-performing model lies towards the top-right of the plot. We use $K=65$k.
@@ 5.1 Implementation Details
We learn representations on the common ImageNet ILSVRC-2012 [9] and its smaller ImageNet-100 subset [32] using a ResNet-50 [15] encoder. All runs of SynCo are based on MoCo-v2 [8]. We developed our approach on top of the official public implementation of MoCo and reproduced it on our setup. We run all experiments on 4 NVIDIA A40 GPUs. For linear classification on ImageNet, we follow the common protocol and report results on the validation set. We report performance after learning linear classifiers for 100 epochs, with an initial learning rate of $30.0$ ($10.0$ for ImageNet-100), a batch size of 256 and a cosine learning rate schedule. For training, unless stated otherwise, we use $K=65$k ($K=16$k for ImageNet-100). For SynCo, we also have a warm-up of 10 epochs, i.e. for the first epochs we do not synthesize hard negatives. We set SynCo’s hyperparameters $\sigma$, $\delta$, and $\eta$ to 0.01. For hard negative generation, we use the top $N=1024$ hardest negatives, with $N_{1}=N_{2}=N_{3}=256$ and $N_{4}=N_{5}=N_{6}=64$. For ImageNet, we report accuracy for a single-crop testing. For object detection on PASCAL VOC [11] we follow [17] and fine-tune a Faster R-CNN [30], R50-C4 on trainval07+12 and test on test2007. For COCO object detection, we follow [17] and fine-tune a Mask R-CNN [16], R50-C4 on train2017 and test on val2017. We use the open-source Detectron2 [40] code and report the common $AP$, $AP_{50}$ and $AP_{75}$ metrics. Similar to [17], we do not perform hyperparameter tuning for the object detection task. Detailed implementation details are provided in the supplementary material.
@@ 5.2 Linear Evaluation on ImageNet
We evaluate the SynCo representation by training a linear classifier on top of the frozen features that were pretrained on ImageNet, following the procedure described in [23, 22, 6, 13, 36]. Full details of the linear evaluation protocol are described in the supplementary material. With a standard ResNet-50 and for 200 epoch pretraining, SynCo obtains $67.9\%\pm 0.16\%$ top-1 accuracy and $88.0\%\pm 0.05\%$ top-5 accuracy (Table 1). Specifically, SynCo achieves a +0.4% top-1 accuracy improvement over MoCo-v2 and a +1.0% improvement over MoCHI with 200 epochs of pretraining, leveraging synthetic hard negatives. While MoCHI, which employs hard negatives, achieves lower performance than MoCo-v2, our method, which generates synthetic hard negatives, not only avoids this drop in performance but actually improves it. SynCo also surpasses the state-of-the-art methods SimCLR and SimSiam. When training for 800 epochs, SynCo obtains 70.6% top-1 accuracy (89.8% top-5 accuracy) (Table 2). This represents a +1.9% improvement in top-1 accuracy over MoCHI and a -0.5% decrease compared to MoCo-v2 under the same 800 epochs pretraining settings. These results show that SynCo achieves higher accuracy than its competitors with only a minor computational overhead for generating synthetic hard negatives, while also enabling faster training convergence.
@@ 5.3 Semi-supervised Training on ImageNet
We evaluate the performance obtained when fine-tuning SynCo representation, pretrained for 800 epochs, on a classification task with a small subset of ImageNet’s train set, this time using label information. We follow the semi-supervised protocol of [23, 45, 6, 13, 36], and use the same fixed splits of respectively 1% and 10% of ImageNet labeled training data as in [6]. Our complete semi-supervised learning setup is detailed in the supplementary material. Despite not being specifically designed for this task, SynCo shows promising results (Table 3). With 1% labeled data, SynCo achieves 50.8% $\pm$ 0.21% top-1 accuracy and 77.5% $\pm$ 0.12% top-5 accuracy. With 10% labeled data, SynCo reaches 66.6% $\pm$ 0.19% top-1 accuracy and 88.0% $\pm$ 0.10% top-5 accuracy. We observe that despite a lower volume of labeled data, SynCo competes well with established semi-supervised methods like SimCLR, BYOL, and Barlow Twins.
@@ 5.4 Transferring to Detection
We evaluate the SynCo representation, pretrained for 200 epochs, by applying it to the detection task. We follow the protocol of [17], and as used in [20] for a fair comparison. Detailed configurations for object detection experiments are provided in in the supplementary material. Table 4 presents the results of object detection on PASCAL VOC and COCO datasets. SynCo displays consistent gains over both the supervised baseline and MoCo-v2. Specifically, SynCo, surpasses MoCo-v2 and performs on par with MoCHI in the PASCAL VOC detection task. On the more challenging COCO dataset, SynCo achieves new state-of-the-art results in both bounding box detection and instance segmentation. Additional results comparing SynCo and MoCo-v2 at different training epochs are presented in the supplementary material.
@@ 6 Conclusion
In this paper, we built upon the observation that harder negatives are needed for more effective self-supervised contrastive learning. Based on that, we presented SynCo, a synthetic hard negative generation approach that improves the quality of representations learned in an unsupervised way, offering better transfer learning performance and utilization of the embedding space. We showed that our method learns generalizable representations, which is important considering the high compute cost of self-supervised learning. We found that multiple SynCo configurations provide considerable gains, and that using synthetic hard negatives consistently improves transfer learning performance.
While our experiments primarily utilized the MoCo framework, the proposed hard negative generation strategies are general and can be applied to any contrastive learning method that benefits from hard negatives, including SimCLR[6], CPC [36], PIRL [25], and others [42, 39, 21, 10, 34]. These methods, which employ the InfoNCE loss function (or its variants [6, 10]) and instance discrimination as the pretext task, can benefit from the enhanced hard negative generation strategies proposed by SynCo. The introduction of synthetic hard negatives can provide these methods with more challenging and informative contrasts, potentially leading to better feature representations. Also, the applicability of SynCo extends beyond visual representation learning. The concept of hard negatives can be applied to various domains such as natural language processing, audio processing, and other areas where contrastive learning is relevant.

### v3

@@ 4.1 Hard Negatives Generation
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\text{sim}(\cdot,\cdot)$ is the similarity function and $\delta$ controls the perturbation magnitude. The perturbed embeddings are then normalized and added to the negative logits. Perturbation modifies the embeddings of the hardest negative features based on the gradient of the similarity function, creating synthetic negatives that are slightly adjusted to be more challenging for the model.
Adversarial hard negatives (type 6). For each query $\mathbf{q}$, we propose to generate $N_{6}$ hard negative features by applying adversarial perturbations to the hardest negative features to maximize their similarity to the query embeddings. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{6}=\{\mathbf{h}_{1}^{6},\mathbf{h}_{2}^{6},\ldots,\mathbf{h}_{N_{6}}^{6}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{6}\in H^{6}$ would be given by:
@@ 4.2 Discussion on Synthetic Hard Negatives
Intuition of synthetic hard negatives. Each SynCo strategy imrpoves model generalization through challenging contrasts. Type 1 interpolates between query and hard negatives, increasing sample diversity throughout training. Type 2 extrapolates beyond the query, pushing representation space boundaries and improving robustness to difficult contrasts. Type 3 combines pairs of hard negatives, encouraging more generalized and robust feature learning. Type 4 injects Gaussian noise, promoting invariance to minor feature fluctuations and enhancing generalization. Type 5 modifies embeddings based on similarity gradients, refining discriminatory power by directing the model towards harder negatives. Type 6 applies adversarial perturbations, creating the most challenging contrasts to distinguish deceptively similar samples. A simplified example illustrating the synthetic hard negative generation is provided in Figure 4. The latter figure visualizes the different types of synthetic hard negatives generated from a query and three original negative features in a 2D representation space. Dotted lines connect the query to extrapolated negatives and between pairs of mixup negatives, illustrating the relationships used to generate these synthetic negative features.
@@
Figure 4: A toy example illustrating the six types of synthetic hard negatives generated by SynCo from a query and three original negative features in a 2D representation space.
@@
Is the proxy task more difficult? Figure 3 depicts the proxy task performance for different configurations of SynCo. We observe that incorporating synthetic negatives leads to faster learning and improved performance. Each type of synthetic negative accelerates learning compared to the MoCo-v2 baseline, with the full SynCo configuration showing the most significant improvement (see Figure 6(a)) and the lowest final proxy task performance. This indicates that SynCo presents the most challenging proxy task. This is evidenced by $\max\ell(\mathbf{h}_{k}^{i})>\max\ell(\mathbf{n}_{j})$, where $\mathbf{h}_{k}^{i}\in H^{i}$ are synthetic negatives and $\mathbf{n}_{j}\in\tilde{\mathcal{Q}}_{N}$ are original negatives. Through SynCo, we modulate proxy task difficulty via synthetic negatives, pushing the model to learn more robust features.
@@
Evaluating the usage of the representation space. To assess learned representations, we employ alignment and uniformity metrics proposed by [44]. These metrics provide insights into representation space utilization, with alignment quantifying the grouping of similar samples and uniformity measuring representation spread across the hypersphere. Figure 5 presents results for various models using features from the ImageNet-100 validation set. Our findings demonstrate that SynCo significantly enhances the uniformity of representations compared to MoCo-v2 and MoCHI, indicating its superior ability to utilize the representation space in the proxy task. Furthermore, the incorporation of synthetic negatives (types 1 to 6) leads to improved alignment. These results suggest that SynCo’s approach to synthetic negative generation and contrastive learning yields more effective and well-distributed feature representations.
@@
Figure 5: Performance comparison of MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of alignment and uniformity metrics. The x-axis and y-axis represent $-\mathcal{L}_{\text{uniform}}$ and $-\mathcal{L}_{\text{align}}$, respectively. The model with the highest performance is located in the upper-right corner of the chart. We use $K=65$k.
@@ 5.1 Implementation Details
We pretrain SynCo on ImageNet ILSVRC-2012 [10] and its smaller ImageNet-100 subset [38] for ablation studies using a ResNet-50 [18] encoder. Our implementation builds upon MoCo-v2 [8]. For training, unless stated otherwise, we use $K=65$k ($K=16$k for ImageNet-100). For SynCo, we also have a warm-up of 10 epochs, i.e. for the first epochs we do not synthesize hard negatives. We set SynCo’s hyperparameters $\sigma$, $\delta$, and $\eta$ to 0.01. For hard negative generation, we use the top $N=1024$ hardest negatives, with $N_{1}=N_{2}=N_{3}=256$ and $N_{4}=N_{5}=N_{6}=64$. For ImageNet linear evaluation, we train a linear classifier on frozen features for 100 epochs, using a batch size of 256 and a cosine learning rate schedule. Initial learning rates are set to 30.0 for ImageNet and 10.0 for ImageNet-100. To evaluate transfer learning, we apply SynCo to object detection tasks. For PASCAL VOC [12], we fine-tune a Faster R-CNN [35] (R50-C4) on trainval07+12 and test on test2007. For COCO [28], we use a Mask R-CNN [19] (R50-C4), fine-tuning on train2017 and evaluating on val2017. We employ Detectron2 [46] and report standard AP metrics, following [20] without additional hyperparameter tuning. Detailed implementation details are provided in the supplementary material.
@@ 5.2 Linear Evaluation on ImageNet
We evaluate the SynCo representation by training a linear classifier on top of the frozen features that were pretrained on ImageNet, following the procedure described in [26, 25, 6, 14, 42] (details in supplementary material). SynCo obtains $67.9\%\pm 0.16\%$ top-1 accuracy and $88.0\%\pm 0.05\%$ top-5 accuracy after 200 epoch pretraining (Table 1). Specifically, SynCo achieves a +0.4% top-1 accuracy improvement over MoCo-v2 and a +1.0% improvement over MoCHI. While MoCHI, which employs hard negatives, achieves lower performance than MoCo-v2, our method, which generates synthetic hard negatives, not only avoids this drop in performance but actually improves it. SynCo also surpasses the state-of-the-art methods SimCLR and SimSiam. When training for 800 epochs, SynCo obtains 70.6% top-1 accuracy (89.8% top-5 accuracy) (Table 2). This represents a +1.9% improvement in top-1 accuracy over MoCHI. These results show that SynCo achieves higher accuracy than its competitors with only a minor computational overhead for generating synthetic hard negatives, while also enabling faster training convergence.
@@ 5.3 Semi-supervised Training on ImageNet
We evaluate SynCo in a semi-supervised setting using 1% and 10% of labeled ImageNet data, following protocols from [26, 52, 6, 14, 42] (details in supplementary material). We use the same data splits as [6]. Results in Table 3 show SynCo’s competitive performance. With 1% labeled data, SynCo achieves $50.8\%\pm 0.21\%$ top-1 accuracy and $77.5\%\pm 0.12\%$ top-5 accuracy. Using 10% labeled data, it reaches $66.6\%\pm 0.19\%$ top-1 accuracy and $88.0\%\pm 0.10\%$ top-5 accuracy. We observe that despite a lower volume of labeled data, SynCo competes well with established semi-supervised methods like SimCLR, BYOL, and Barlow Twins.
@@ 5.4 Transferring to Detection
We evaluate the SynCo representation, pretrained for 200 epochs, by applying it to the detection task. We follow the protocol of [20]. Detailed configurations for object detection experiments are provided in in the supplementary material. Table 4 presents the results of object detection on PASCAL VOC and COCO datasets. SynCo displays consistent gains over both the supervised baseline and MoCo-v2. Specifically, SynCo, surpasses MoCo-v2 and performs on par with MoCHI in the PASCAL VOC detection task. On the more challenging COCO dataset, SynCo achieves new state-of-the-art results in both bounding box detection and instance segmentation. Additional results comparing SynCo and MoCo-v2 at different training epochs are presented in the supplementary material.
@@ 6 Conclusion
This paper introduces SynCo, a novel approach to contrastive learning that leverages synthetic hard negatives to enhance visual representation learning. By generating diverse and challenging negative samples on-the-fly, SynCo addresses the limitations of existing methods in maintaining a pool of effective hard negatives throughout training. Our comprehensive experiments demonstrate that SynCo not only accelerates the learning process but also leads to more robust and transferable representations. The effectiveness of SynCo is evidenced by its superior performance across various benchmarks, including linear evaluation on ImageNet, semi-supervised learning tasks, and transfer learning to object detection on PASCAL VOC and COCO datasets.
While our experiments primarily utilized the MoCo framework, the proposed hard negative generation strategies are general and can be applied to any contrastive learning method that benefits from hard negatives, including SimCLR[6], CPC [42], PIRL [30], and others [49, 45, 24, 11, 40]. These methods, which employ the InfoNCE loss function (or its variants [6, 11]) and instance discrimination as the pretext task, can benefit from the enhanced hard negative generation strategies proposed by SynCo. The introduction of synthetic hard negatives can provide these methods with more challenging, informative contrasts, potentially leading to better feature representations. Also, SynCo’s applicability extends beyond visual representation learning. The concept of hard negatives can be applied to various domains such as natural language processing, audio processing, and other areas where contrastive learning is relevant.
Broader Impact
The presented research should be categorized as research in the field of unsupervised learning. This work may inspire new algorithms, theoretical, and experimental investigation. The algorithm presented here can be used for many different vision applications and a particular use may have both positive or negative impacts, which is known as the dual use problem. Besides, as vision datasets could be biased, the representation learned by SynCo could be susceptible to replicate these biases.